# LLM4GEN: Leveraging Semantic Representation of LLMs for Text-to-Image Generation

### v1

@@ 4.1. DensePrompts Benchmark
4.2. Training Dataset Construction
Current text-to-image diffusion models use text-image pairs collected from the Internet, where the text descriptions are often brief and disconnected. This leads to a weak correlation between images and text, limited dense semantic information, and necessitates a large number of image-text pairs for model convergence.
In this paper, we propose a 1M LAION-refined dataset, including the refined image descriptions of the open-source LAION (Schuhmann et al., 2021) dataset. We select 1M images from LAION (Schuhmann et al., 2021) with an aesthetic score (Schuhmann, [n. d.]) exceeding 6.5, a minimum short edge resolution of 512 pixels, and a maximum aspect ratio of 1.5. We utilize LLaVA-7B (Liu et al., 2023) to generate dense and highly descriptive captions for each image. The process of dataset construction is illustrated in Fig. 4 (b).
5. Experiments
Figure 6. Aesthetic Score and CLIP Score (%) on DensePrompts benchmark.
5.1. Experimental Details
Framework. In this paper, we explore LLM4GEN based on SD1.5 (Rombach et al., 2022) and SDXL (Podell et al., 2023), denoted as LLM4GENSD1.5 and LLM4GENSDXL. We utilize T5-XL (Raffel et al., 2020) and CLIP text encoder as the text tower. The sequence length of the LLMs is set to 128, enhancing the representation of conditional text embedding.
Implementation Details. We employ AdamW (Loshchilov and Hutter, 2017) optimizer to train our models. The learning rates are set to 2e-5 and 1e-5 for LLM4GENSD1.5 and LLM4GENSDXL, respectively. The batch size is set to 256. The training steps are set to 20k and 40k. Additionally, we further train LLM4GENSDXL using 20K high-quality data with 1024 resolution. The final LLM4GENSD1.5 model is trained on 8 80G A100 for 2 days while 4 days for LLM4GENSDXL model. During inference, we utilize DDIM sampler (Song et al., 2020a) for sampling, setting the number of time steps to 50 and the classifier free guidance scale to 7.5.
Evaluation Benchmarks. We comprehensively evaluate proposed LLM4GEN via four primary benchmarks:
  1. (1)
    MSCOCO Dataset (Lin et al., 2014) We randomly select 30k images from MSCOCO dataset and assess both sample quality and image-text alignment of generated images. The Fréchet Inception Distance (FID) (Heusel et al., 2017), Inception Score (IS) (Salimans et al., 2016), and CLIP Score (Radford et al., 2021) are used for evaluation.
  2. (2)
    T2I-CompBench (Huang et al., 2023) We employ various compositional prompts to assess textual attributes, including aspects such as color, shape, and texture, as well as attribute binding.
  3. (3)
    DensePrompts Our proposed benchmark involves extensive textual descriptions comprising over 7,000 dense prompts. The CLIP Score and Aesthetic Score are used for evaluation.
  4. (4)
    User Study We randomly select 100 prompts from our proposed DensePromtps benchmark and 100 prompts from T2I-CompBench. Subsequently, we enlist 20 participants for the user study.
5.2. Performance Comparisons and Analysis
5.2.1. Fidelity assessment on MSCOCO benchmark
Experimental results on MSCOCO benchmark are shown in Tab. 2. LLM4GEN notably enhances the sample quality and image-text alignment, resulting in improvements of 1.79 and 0.54 on FID compared to SD1.5 and SDXL, respectively. Furthermore, we assess the performance of SD1.5 after extensive fine-tuning with the same training dataset. This modified version, SD1.5 (ft), surpasses the original SD1.5, yet LLM4GENSD1.5 still exhibits superior performance over SD1.5 (ft). This underscores the potent representation of our proposed LLM4GEN and its contribution to text-to-image generation.
@@ 5.2.1. Fidelity assessment on MSCOCO benchmark
Figure 7. Results on user study regarding the sample quality and image-text alignment of different models.
5.2.2. Evaluation on T2I-CompBench
For T2I-CompBench comparison, we select the recent text-to-image generative models for comparison, e.g., Composable Diffusion (Liu et al., 2022), Structured Diffusion (Feng et al., 2022), Attn-Exct v2 (Chefer et al., 2023), GORS (Huang et al., 2023), DALLE 2 (Ramesh et al., 2022), PixArt-$\alpha$ (Chen et al., 2023b), ELLASDXL (Hu et al., 2024), SD1.5(Rombach et al., 2022), and SDXL(Podell et al., 2023). Experimental results shown in Tab. 1 demonstrate the distinctive performance of LLM4GENSDXL in T2I-CompBench evaluation, underlining its advancements in attribute binding, object relationship, and mastery in rendering complex compositions. LLM4GEN shows considerable improvement in color, shape, and texture, showcasing enhancements up to +9.60% in color, +11.52% in texture, and +3.26% in shape with SDXL, respectively. LLM4GENSDXL also marks considerable progress in both spatial and non-spatial evaluations, with 2.27% and 0.84% lift, respectively. Furthermore, when compared with the recent PixArt-$\alpha$, which employs T5-XL as its text encoder, LLM4GENSDXL surpasses it in several aspects, such as a notable 3.43% lead in color metric. Moreover, LLM4GENSDXL outperforms ELLASDXL. These results verify the potent synergy of LLMs representations in augmenting the sample quality and image-text alignment of diffusion models.
@@ 5.2.2. Evaluation on T2I-CompBench
Figure 8. A comparative analysis of LLM4GEN and other state-of-the-art diffusion models using PartiPrompts (Yu et al., 2022) and our proposed DensePrompts as prompts. The last row represents the prompts used.
5.2.3. Evaluation on DensePrompts
[…]
@@ 5.2.3. Evaluation on DensePrompts
To thoroughly evaluate our proposed LLM4GEN framework, we present the qualitative results on the short prompts provided by PartiPrompts (Yu et al., 2022) in the first 4 columns and on the dense prompts provided by DensePrompts in the last 3 columns in Fig. 8. The results indicate that our proposed LLM4GENSD1.5 and LLM4GENSDXL exhibit strong text-image alignment and superior dense prompt generation compared to the recent PixArt-$\alpha$ and Playground v2, especially in handling the multiple objects and attribute binding.
5.2.4. User Study
@@
We conduct the user study on various combinations of existing methods and LLM4GENSDXL. For each pairing, we assess two criteria: sample quality and image-text alignment. Users are tasked with evaluating the aesthetic appeal and semantic understanding of images with identical text to determine the superior one based on these assessment criteria. Subsequently, we compute the percentage scores for each model, as shown in Fig. 7. The results showcase our LLM4GENSDXL exhibits comparative advantages over both SD1.5 and SDXL. Specifically, LLM4GENSDXL achieves 60.3% and 66.5% higher voting preferences compared to SDXL in terms of Aesthetic and Semantic, respectively. Notably, LLM4GENSDXL also delivers competitive results when compared to DALL-E 3.
Table 3. Impact (%) of the designed Cross-Adapter Module. The symbol ✗ denotes cases where the model fails to generate images without the original text encoder.
@@
(4)  CrossAttention + Concat 45.34 43.28 51.52
5.3. Ablation Studies
Table 4. Impact (%) of Different LLMs based on SD1.5.
LLMs Attribute Binding
Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$
SD1.5 (Rombach et al., 2022) 37.65 35.76 41.56
Llama-2/7B (Touvron et al., 2023) 43.21 40.12 48.91
Llama-2/13B (Touvron et al., 2023) 43.98 41.03 49.21
T5-XL (Raffel et al., 2020) 45.34 43.28 51.52
Impact of Cross-Adapter Module. Due to limited computing sources, we evaluate the impact of various architectural enhancements on SD1.5, as outlined in Tab. 3. Our configurations explore different methods for integrating LLMs embeddings: (1) the baseline SD1.5 model, (2) MLP or CrossAttention, which utilizes a simple linear layer or cross-attention layer to transform LLM embeddings, (3) MLP + Concat, representing a process where LLMs embeddings are projected to the same dimension as the original text embeddings before concatenation, and (4) CrossAttention + Concat, our innovative approach detailed in Sec. 3.2. Results show that configuration (2) is incapable of generating images from text, likely due to a misalignment between LLMs and the latent vector, necessitating substantial resources for alignment, as mentioned in (Hu et al., 2024) and (Wu et al., 2023b). Interestingly, simply concatenating the original text embeddings (configuration 3) provides a significant boost over base SD1.5, with a 2.59% improvement in color. This suggests that direct representation alignment between LLMs and the latent vector is challenging, and enhancing the original text embeddings with LLM embeddings is sufficient to improve image-text alignment. Furthermore, the result of our carefully crafted Cross-Adapter Module (configuration 4) achieves a 5.10% increase in color over configuration 3. This emphasizes the substantial benefits of incorporating our Cross-Adapter Module to enrich the representation of the original text encoder and the image-text alignment of generated images.
Table 5. Training resources comparison, including the scale of training data (#Images) and computing cost (#GPU days). The performance (%) is evaluated on the Color in T2I-CompBench.
@@ 5.3. Ablation Studies
PixArt-$\alpha$ (Chen et al., 2023b) 25 753 68.86
ParaDiffusion (Wu et al., 2023b) 500 392 -
ELLASDXL (Hu et al., 2024) 30 112 72.60
LLM4GENSDXL 2 32 73.29
Figure 9. Performance metrics of LLM4GEN based on different LLMs across $\lambda$ values from 0 to 2.
Impact of Different LLMs. The analysis encompasses a comparative evaluation between base SD1.5 and the enhancements achieved through the integration of Llama-2/7B, Llama-2/13B, and T5-XL. As depicted in Tab. 4, the inclusion of any LLM improves upon the performance of SD1.5. Importantly, Llama-v2/13B outperforms Llama-v2/7B, demonstrating that LLMs with greater capacity excel in extracting more nuanced semantic embeddings. Furthermore, when compared to decoder-only LLMs, T5-XL encoder demonstrates advantages in semantic comprehension, confirming its superior suitability for enhancing text-to-image generation.
Impact of hyperparameter $\lambda$. As depicted in Eq. 7, the hyperparameter $\lambda$ is used to regulate the weight of the LLM’s embedding injected into the original text embedding. We evaluate the impact of $\lambda$ in Fig. 9 with FID and CLIP Score for LLM4GENSD1.5 on MSCOCO dataset across different LLMs. The $\lambda$ varies from 0 to 2 in 0.2 increments. As the $\lambda$ increases, we observe an initial enhancement in the model’s performance, followed by a slight decline. This pattern indicates that integrating LLMs representation into that of the original text encoder can consistently improve the image-text alignment beyond the capabilities of the original SD1.5, highlighting the beneficial impact of LLMs on semantic enrichment. However, setting higher values of $\lambda$ does not lead to optimal performance, which we suspect stems from a misalignment between LLMs representation and the diffusion model. The best performance is achieved when $\lambda$ is set to 1.0.
5.4. Further Analysis
Training Efficiency. When evaluating the effectiveness of integrating LLMs into text-to-image generation models, LLM4GENSDXL stands out for its remarkable efficiency and performance. LLM4GEN achieves significant reductions in both training data requirements and computational costs. It utilizes only 2M data, a 90% reduction compared to ELLA (Hu et al., 2024), and demands merely 32 GPU days for training, drastically lower than PixArt-$\alpha$ (25 million data, 753 GPU days) and ParaDiffusion (500 million data, 392 GPU days). Despite this, LLM4GENSDXL achieves a superior color metric performance of 73.29%. This notable difference underscores LLM4GEN’s ability to substantially reduce both training data and computational costs while establishing a new standard for performance efficiency.
@@ 5.4. Further Analysis
Figure 10. Cross-attention visualization (Tang et al., 2022) for two generated images. The two rows are SD1.5 and LLM4GENSD1.5, respectively.
@@
Cross-attention Visualization. Fig. 10 demonstrates the cross-attention visualization (Tang et al., 2022) of SD1.5 and LLM4GENSD1.5, respectively. The heatmaps indicate that our proposed LLM4GEN method exhibits a greater capacity to capture relationships between attributes, such as ”blue” and ”sheep”, as illustrated in Fig. 10 (a). We attribute this capability to the enhanced semantic richness facilitated by the robust representation of LLMs.
More visualization and experimental results are shown in the Supplementary Materials.
6. Conclusion
@@
In this paper, we propose LLM4GEN, an end-to-end text-to-image generation framework. Specifically, we design an efficient Cross-Adapter Module to leverage the powerful representation of LLMs to enhance the original text representation of diffusion models. Despite using fewer training data and computational resources, LLM4GEN outperforms current state-of-the-art text-to-image diffusion models in sample quality and image-text alignment. Additionally, we introduce LAION-refined dataset and DensePrompts benchmark, which promote generating images with dense information and establish a comprehensive evaluation, respectively. Furthermore, we aim to explore the potential of LLM4GEN within the transformer-based diffusion model. We hope this work will pave the way for new research directions in text-to-image generation and provide insights into how LLMs can contribute to improving the performance of text-to-image generation models.
Supplementary Materials: LLM4GEN: Leveraging Semantic Representation of LLMs for Text-to-Image Generation
7. Additional Analysis
We provide the additional analysis and experimental results on our proposed LLM4GEN in the Supplementary Materials.
Formualtion Derivation We provide a formula derivation of Eq. (8):
(11) $x$ $=\operatorname{CA}(x,c_{t}^{\prime})$
$=\operatorname{softmax}\left(Q^{\prime}\cdot{K^{\prime}}^{T}\right)\cdot V^{\prime}$
$=\operatorname{softmax}\left(W_{Q}^{\prime}(x)\cdot{W_{K}^{\prime}(c_{t}^{\prime})}^{T}\right)\cdot W_{V}^{\prime}(c_{t}^{\prime})$
$=\operatorname{softmax}\left(W_{Q}^{\prime}(x)\cdot{W_{K}^{\prime}([\lambda\cdot c_{l}^{\prime},c_{t}])}^{T}\right)\cdot W_{V}^{\prime}([\lambda\cdot c_{l}^{\prime},c_{t}])$
$=\lambda\cdot\operatorname{softmax}\left(W_{Q}^{\prime}(x)\cdot{W_{K}^{\prime}(c_{l}^{\prime})}^{T}\right)\cdot W_{V}^{\prime}(c_{l}^{\prime})$
$\quad+\operatorname{softmax}\left(W_{Q}^{\prime}(x)\cdot{W_{K}^{\prime}(c_{t})}^{T}\right)\cdot W_{V}^{\prime}(c_{t})$
$=\lambda\cdot\operatorname{CA}(x,c_{l})+\operatorname{CA}(x,c_{t})$
where $x$ denotes the latent noise, CA is the cross-attention module within the UNet module, which receives $z$ as the query and $c_{t}^{\prime}$ as the key and value. And the $W_{Q}^{\prime},W_{K}^{\prime},W_{V}^{\prime}$ is the projector in the UNet. In this manner, the concatenation operation in Eq. (7) is equal to fuse the LLM-guided semantic feature to the latent noise for better text-to-image alignment.
Algorithm. The proposed LLM4GEN is further illustrated in Algorithm 2.
Algorithm 2 LLM4GEN Pipeline
1:  Input: Pretrained UNet $\epsilon$, pre-trained text encoder $T_{\phi}$, pre-trained LLM $T_{L}$, Cross-Adapter Module $M$, LLaVA-7B $A$, training image-text pairs $\mathbb{S}=\{\mathbb{I},\mathbb{P}\}$
2:  Offline Process: Apply $A$ to enrich the captions of $\mathrm{S}$, get $\hat{\mathbb{P}}$ = $A$($\mathrm{I}$) to replace the original $\mathbb{P}$.
3:  $\rightarrow$ Begin Training.
4:  Freeze $T_{\phi}$ and $T_{L}$, training $M$ and $\epsilon$.
5:  while ${L_{simple}}$ not converged do
6:     Sample $p$ from $\mathbb{P}$; $t=T$
7:     ${c_{t}^{\prime}}=M(T_{\phi}(p),T_{L}(p))$ using Eq. (7).
8:     Calculate ${L_{simple}}$ using Eq. (9).
9:     Backward ${L_{simple}}$ and update $\epsilon$, $M$.
10:  end while
11:  $\rightarrow$ End Training.
Figure 11. Comparison with ParaDiffusion (Wu et al., 2023b). The generation images of ParaDiffusion are from the original paper.
Figure 12. Comparison with ELLA (Hu et al., 2024). The generation images of ParaDiffusion are from the original paper.
8. Additional Experimental Results
Comparison with ParaDiffusion based on dense prompts. To further compare the proposed LLM4GEN with ParamDiffusion, we take three prompts provided in (Wu et al., 2023b), and generate images using our LLM4GENSD1.5 and LLM4GENSDXL. We select the generation images of ParaDiffusion from the original literature (Wu et al., 2023b). The results demonstrate that our proposed LLM4GEN can generate semantic-alignment images based on long textual descriptions and alleviate bad cases in ParaDiffusion, such as confusion between character generation and background in Fig. 11 (c). LLM4GENSD1.5 even can generate high-quality and dense prompts alignment images than SDXL(Podell et al., 2023), such as Fig. 11 (b). It demonstrates our proposed methods can efficiently integrate strong semantic representations of LLMs into text-to-image diffusion models to enhance image-text alignment.
Comparison with ELLA (Hu et al., 2024). ELLA (Hu et al., 2024) also utilizes LLMs for text-to-image generation, with the aim of aligning LLMs with diffusion models from scratch, incurring significant training costs. In contrast to ELLA, our proposed LLM4GEN enhances the original text encoder directly with robust LLM semantic embeddings. This approach benefits from the powerful capabilities of LLM without high-cost training and computing sources. Compared to ELLASDXL, our model requires only 10% of the training data, but still produces high-quality images using the prompts provided in (Hu et al., 2024). The visualization comparison is presented in Fig. 12.
Figure 13. LLMGENSD1.5 can be integrated with ControlNet (Zhang et al., 2023) to generate images with guided poses.
Compatible with ControlNet (Zhang et al., 2023). We apply LLM4GEN with ControlNet (Zhang et al., 2023), as shown in Fig. 13. We can see that LLM4GENSD1.5 can be integrated with exsiting guided tools like ControlNet. LLM4GENSD1.5 can be compatible with these standard methods while generates more consistent and text-align images.

### v2

@@ Introduction
Table 2: Quantitative comparison on text-to-image generation models on the subset of MSCOCO (Lin et al. 2014) dataset.
Figure 5: Aesthetic Score and CLIP Score (%) on DensePrompts benchmark.
@@
Figure 6: Results on user study regarding the sample quality and image-text alignment of different models.
@@
Figure 7: A comparative analysis of LLM4GEN and other state-of-the-art diffusion models using PartiPrompts (Yu et al. 2022) and our proposed DensePrompts as prompts. The last row represents the prompts used.
Performance Comparisons and Analysis
Fidelity assessment on MSCOCO benchmark Experimental results on MSCOCO benchmark are shown in Tab. 2. LLM4GEN notably enhances the sample quality and image-text alignment, resulting in improvements of 1.79 and 0.54 on FID compared to SD1.5 and SDXL, respectively. Furthermore, we assess the performance of SD1.5 after extensive fine-tuning with the same training dataset. This modified version, SD1.5 (ft), surpasses the original SD1.5, yet LLM4GENSD1.5 still exhibits superior performance over SD1.5 (ft). This underscores the potent representation of our proposed LLM4GEN and its contribution to text-to-image generation.
Evaluation on T2I-CompBench For T2I-CompBench comparison, we select the recent text-to-image generative models for comparison, e.g., Composable Diffusion, Structured Diffusion, Attn-Exct v2, GORS, DALLE 2, PixArt-$\alpha$, ELLASDXL, SD1.5, and SDXL. Experimental results shown in Tab. 1 demonstrate the distinctive performance of LLM4GENSDXL in T2I-CompBench evaluation, underlining its advancements in attribute binding, object relationship, and mastery in rendering complex compositions. LLM4GEN shows considerable improvement in color, shape, and texture, showcasing enhancements up to +12.90% in color, +5.16% in shape, and +14.49% in texture with SDXL, respectively. LLM4GENSDXL also marks considerable progress in both spatial and non-spatial evaluations, with 3.80% and 1.00% lift, respectively. Furthermore, when compared with PixArt-$\alpha$, which employs T5-XL as its text encoder, LLM4GENSDXL surpasses it in several aspects, such as a notable 7.73% lead in color metric. Moreover, LLM4GENSDXL outperforms ELLASDXL. These results verify the potent synergy of LLMs representations in augmenting the sample quality and image-text alignment of diffusion models.
Evaluation on DensePrompts We compare our LLM4GEN with PixArt-$\alpha$, Playground v2, SD1.5, and SDXL on our DensePrompts benchmark. As shown in Fig. 5, LLM4GENSDXL achieves the highest Aesthetic Score and CLIP Score among these models. PixArt-$\alpha$ outperforms SDXL due to its T5-XL text encoder for dense prompts. LLM4GEN excels in understanding and interpreting dense prompts, resulting in high-quality images with strong image-text alignment. This performance is attributed to the powerful representation of LLMs and the effective adaptation of the original CLIP text encoder via our CrossAdapter Module. We attribute this performance to th powerful representation of LLMs and the effective adaptation of the original CLIP text encoder via our CrossAdapter Module.
@@
Quantitive Results. To thoroughly evaluate our proposed LLM4GEN framework, we present the qualitative results on the short prompts provided by PartiPrompts (Yu et al. 2022) in the first 4 columns and on the dense prompts provided by DensePrompts in the last 3 columns in Fig. 7. The results indicate that our proposed LLM4GENSD1.5 and LLM4GENSDXL exhibit strong text-image alignment and superior dense prompt generation compared to the recent PixArt-$\alpha$, especially in handling the multiple objects and attribute binding.
@@
User Study We conduct the user study on various combinations of existing methods and LLM4GENSDXL. For each pairing, we assess two criteria: sample quality and image-text alignment. Users are tasked with evaluating the aesthetic appeal and semantic understanding of images with identical text to determine the superior one based on these assessment criteria. Subsequently, we compute the percentage scores for each model, as shown in Fig. 6. The results showcase our LLM4GENSDXL exhibits comparative advantages over both SD1.5 and SDXL. Specifically, LLM4GENSDXL achieves 60.3% and 66.5% higher voting preferences compared to SDXL in terms of Aesthetic and Semantic, respectively. Notably, LLM4GENSDXL also delivers competitive results when compared to DALL-E 3.
Ablation Studies
Impact of Cross-Adapter Module
Due to limited computing sources, we evaluate the impact of various architectural enhancements on SD1.5, as outlined in Tab. 4. Our configurations explore different methods for integrating LLMs embeddings: (1) the baseline SD1.5 model, (2) SD1.5 finetune-CLIP, the result of fine-tuning the original text-encoder of SD1.5, (3) MLP or CrossAttention, which utilizes a simple linear layer or cross-attention layer to transform LLM embeddings, (4) MLP + Concat, representing a process where LLMs embeddings are projected to the same dimension as the original text embeddings before concatenation, (5) CrossAttention + Concat, (6) CLIP as Q and LLM as KV, refering to converting the position of Q and KV in CAM. Results show that configuration (2) indeed brings an improvement of the original SD1.5, yet, due to the limited semantic representation of CLIP, the results still remain subpar. Interestingly, simply concatenating the original text embeddings (configuration 3 & 4) provides a significant boost over base SD1.5. This suggests that direct representation alignment between LLMs and the latent vector is challenging, and enhancing the original text embeddings with LLM embeddings is sufficient to improve image-text alignment. In our LLM4GEN, the LLM representation is employed as Q, while the original text encoder serves as K and V. We also examine the impact of rearranging the position of Q and KV in the CAM module. The results, as demonstrated in configuration 6, indicate that our LLM4GEN (configuration 7) exceeds it, showcasing a 3.65% enhancement in color. This emphasizes the substantial benefits of incorporating our Cross-Adapter Module to enrich the representation of the original text encoder and the image-text alignment of generated images.
LLMs Attribute Binding
Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$
SD1.5 (Rombach et al. 2022) 37.65 35.76 41.56
Llama-2/7B (Touvron et al. 2023) 43.21 40.12 48.91
Llama-2/13B (Touvron et al. 2023) 44.98 41.03 49.21
T5-XL (Raffel et al. 2020) 47.34 45.35 55.39
Table 3: Impact (%) of Different LLMs based on SD1.5.
@@
(5)  CrossAttention + Concat 42.31 39.43 46.21
(6)  CLIP as Q and LLM as KV 43.69 42.82 49.56
(7)  Ours 47.34 45.35 55.39
Table 4: Impact (%) of the designed Cross-Adapter Module.
Impact of Different LLMs
The analysis encompasses a comparative evaluation between base SD1.5 and the enhancements achieved through the integration of Llama-2/7B, Llama-2/13B, and T5-XL. As depicted in Tab. 3, the inclusion of any LLM improves upon the performance of SD1.5. Importantly, Llama-v2/13B outperforms Llama-v2/7B, demonstrating that LLMs with greater capacity excel in extracting more nuanced semantic embeddings. Furthermore, when compared to decoder-only LLMs, T5-XL encoder demonstrates advantages in semantic comprehension, confirming its superior suitability for enhancing text-to-image generation.
Further Analysis
@@
PixArt-$\alpha$ (Chen et al. 2023b) 25 753 68.86
ParaDiffusion (Wu et al. 2023) 500 392 -
ELLASDXL (Hu et al. 2024) 30 112 72.60
LLM4GENSDXL 10 50 76.59
Table 5: Training resources comparison, including the scale of training data and computing cost.
Figure 8: Training data scaling analysis.
@@
Figure 9: Cross-attention visualization (Tang et al. 2022) for two generated images. The two rows are SD1.5 and LLM4GENSD1.5, respectively.
Scaling Analysis
As illustrated in Fig. 8, we conduct an extensive analysis of the scalability of our proposed LLM4GEN model with respect to training data. The results conclusively demonstrate that as the quantity of training data increases, the performance of our model exhibits consistent and significant growth, thereby confirming its scalability. However, increasing the dataset scale from 5M to 10M resulted in minimal performance improvement on the generated images. Consequently, we use 10M text-image pairs for training our LLM4GEN.
Training Efficiency
When evaluating the effectiveness of integrating LLMs into text-to-image generation models, LLM4GENSDXL stands out for its remarkable efficiency and performance. LLM4GEN achieves significant reductions in both training data requirements and computational costs. It utilizes only 10 million data, a 66% reduction compared to ELLA, and demands merely 50 GPU days for training, drastically lower than PixArt-$\alpha$ (25 million data, 753 GPU days) and ParaDiffusion (500 million data, 392 GPU days). Despite this, LLM4GENSDXL achieves a superior color metric performance of 73.29%. This notable difference underscores LLM4GEN’s ability to substantially reduce both training data and computational costs while establishing a new standard for performance efficiency.
Cross-attention Visualization.
@@
Fig. 9 shows the cross-attention visualization of SD1.5 and LLM4GENSD1.5, respectively. The heatmaps reveal that our proposed LLM4GEN method demonstrates a superior ability to capture relationships between attributes, such as ”blue” and ”sheep,” as illustrated in Fig. 9 (a). We attribute this enhanced capability to the increased semantic richness afforded by the robust representations of LLMs.
More visualization and experimental results are shown in the Appendix.
Conclusion
@@
In this paper, we propose LLM4GEN, an end-to-end text-to-image generation framework. Specifically, we design an efficient Cross-Adapter Module to leverage the powerful representation of LLMs, thereby enhancing the original text representation of diffusion models. Despite using fewer training data and computational resources, LLM4GEN outperforms current state-of-the-art text-to-image diffusion models in sample quality and image-text alignment. To optimize consistency in entity-attribute relationships of generated images, we design an entity-guided regularization loss. Additionally, we introduce the DensePrompts benchmark to promote the generation of images with dense information and provide a comprehensive evaluation framework. Extensive experiments have shown that our proposed method achieves competitive performance.